### v1

# Generalizable Error Modeling for Search Relevance Data Annotation Tasks

Abstract
Human data annotation is critical in shaping the quality of machine learning (ML) and artificial intelligence (AI) systems. One significant challenge in this context is posed by annotation errors, as their effects can degrade the performance of ML models. This paper presents a predictive error model trained to detect potential errors in search relevance annotation tasks for three industry-scale ML applications (music streaming, video streaming, and mobile apps) and assesses its potential to enhance the quality and efficiency of the data annotation process. Drawing on real-world data from an extensive search relevance annotation program, we illustrate that errors can be predicted with moderate model performance (AUC=0.65-0.75) and that model performance generalizes well across applications (i.e., a global, task-agnostic model performs on par with task-specific models). We present model explainability analyses to identify which types of features are the main drivers of predictive performance. Additionally, we demonstrate the usefulness of the model in the context of auditing, where prioritizing tasks with high predicted error probabilities considerably increases the amount of corrected annotation errors (e.g., 40% efficiency gains for the music streaming application). These results underscore that automated error detection models can yield considerable improvements in the efficiency and quality of data annotation processes. Thus, our findings reveal critical insights into effective error management in the data annotation process, thereby contributing to the broader field of human-in-the-loop ML.
Keywords Error Modeling  $\cdot$ Data Annotation  $\cdot$ Machine Learning  $\cdot$ Data Quality  $\cdot$ Human-in-the-loop ML
1 Introduction
One of the key components of modern machine learning (ML) and artificial intelligence (AI) systems is the ever-increasing volume of human-annotated data used to train, evaluate, monitor, and update such models. The importance of this subject can hardly be overstated, as the quality of ML models is linked to the quality of the data they are trained and evaluated on [1, 2, 3, 4, 5]. With the data annotation market reaching an estimated \$1 billion in 2023 [6], it is important to note that the impact of this industry extends beyond this figure. The revenue that is tied to the ML products relying on human annotations is far greater than the data annotation market itself, underscoring the economic influence and strategic relevance of human annotations. However, a significant challenge lies in the management of annotation errors and their downstream consequences. The effects of these errors can cause setbacks in ML applications, making their management a critical task in maintaining the overall quality and effectiveness of ML models [1].
In supervised ML, data annotation errors can impact model quality in at least two ways: First, mislabeling of training data is one of the main sources of prediction errors, as it prevents models from picking up on the true target classification function [3, 4]. Second, mislabeling of testing data distorts model evaluations, therefore affecting model selection and leading to the deployment of suboptimal models [5]. For example, correcting labeling errors in the widely used ImageNet [7] benchmark dataset would change the rank order of model performance in favor of lower capacity models [5].
Several potential solutions have been proposed to tackle annotation errors. These include developing clear annotation guidelines paired with extensive annotator training [8], generating crowd-sourced annotations and monitoring inter-annotator agreement [9], implementing audits, quality checks, or relabeling by subject matter experts [2], as well as automated error detection techniques [10, 2]. Generally, these practices should be implemented in a way that allows for feedback loops and iterative refinement. For example, detected errors can be used as an opportunity to refine guidelines and provide more training to annotators. Similarly, annotators should be able to ask questions or raise concerns about the annotation process or specific items to help identify problems early [11, 12].
The present paper explores automated error models and their potential to improve the efficiency of the data annotation process, specifically with respect to the quality and turnaround of annotations. We further analyze how automated error detection models can improve related processes such as auditing and relabeling. Compared to previous work addressing annotation errors, our approach introduces a unique emphasis on task-agnostic error prediction. Unlike many prior models, which focus on specific tasks and require separate training for each labeling task, our models are trained to predict errors directly from a combination of task features and annotator features, offering a higher degree of generalizability.
1.1 Background and Related Work
Automated error detection techniques aim to identify instances that may be labeled incorrectly or inconsistently within a given dataset [10]. Once these potential errors are detected, they can be forwarded for manual review by human annotators or processed via annotation error correction methods. Error detection has been implemented across diverse domains in natural language processing [13, 14, 10], and image classification [5, 2]. However, the logic of error modeling can be applied to a variety of supervised ML tasks.
One key distinction between error detection techniques is whether the approach is variation-based, rule-based, or model-based [10]. Variation-based approaches generally rely on inputs labeled by multiple annotators. For example, in settings where only one label can be correct, disagreement across annotators can indicate annotation errors. Similarly, annotator disagreement can be captured by computing variation metrics like label entropy across annotators, such that high entropy would be indicative of annotation errors [15]. While variation-based error detection methods are easy to implement, one major disadvantage is that they are limited to crowd-sourced scenarios, where more than one annotator label is produced for each instance. Requiring multiple labels can introduce significant costs to annotation processes.
Rule-based techniques, on the other hand, are useful when there are clear annotation guidelines that can be translated into formal rules. In such situations, annotations that are impossible given the set of rules would be flagged as incorrect. For example, errors can be flagged if an annotator picks a label that is not applicable to the given task. In language annotation tasks, n-grams that are lexically or morphologically impossible would fall into this category [16]. However, while rule-based approaches can prove to be highly effective, they are typically tailored to a very specific task, imposing limits on generalizability. Additionally, they only apply to a specific subset of errors, namely those that are the result of objective rule violations.
Model-based error detection approaches typically utilize probabilistic classifiers designed to detect annotation errors. Past work has frequently relied on the models predicting an annotation label (typically a multi-class classification problem) and subsequently comparing the predicted label to a label provided by a human annotator. If there is disagreement between the human annotation and the model-generated annotation, the annotation can be marked as inconsistent and additionally directed to a human auditor [17, 18]. Alternatively, model confidence scores or predicted probabilities can provide an indication of task difficulty [19]. Implicitly, these approaches rely on the assumption that models perform worse on instances that are more difficult for human annotators as well. Consequently, instances that show low model performance - either by deviation from human annotations or by low model confidence scores - would be more likely to contain annotation errors.
However, it is important to note that a model’s confidence score on the labeling task itself may not always be a reliable indicator of a human annotator’s potential error. For instance, low model confidence could be due to the model’s limitations in representing complex, nuanced, or context-specific tasks rather than indicating an annotation error. Conversely, there may be cases where a model is highly confident and yet wrong due to biases in training data or overly simplistic decision boundaries. Therefore, while using model confidence scores can certainly be a useful tool in identifying potential annotation errors, it should be treated as part of a broader toolkit and used in conjunction with other methods. For a more comprehensive approach to error detection, it might be beneficial also to consider other factors such as the context of the task, the difficulty level of the task, and the reliability and expertise of the human annotators themselves. Relatedly, this approach can be less effective in identifying systematic errors, where both the model and the human annotators consistently misinterpret certain types of instances. Such errors could go undetected as model predictions and human annotations would agree, despite being incorrect. Finally, predicting labels may suffer from low generalizability, as a separate model needs to be trained for each specific labeling task, which can be computationally expensive and may not be feasible for all tasks, especially those requiring complex models.
As an alternative approach explored in this paper, a model can be used to directly predict whether a human label is correct or not (typically a binary classification problem) without explicitly predicting the underlying label itself. The advantage of this approach is that it is often easier to predict the presence of an error compared to predicting the correct label. Relatedly, directly predicting the correctness of a label can be computationally more efficient, as it does not necessarily need to consider all possible label classes, which can be vast in certain tasks. Overall this approach makes it easier to train, evaluate, and maintain such models even when training data are limited. Additionally, error detection models can be set up to be task agnostic: a model predicting errors instead of labels may generalize to a wider range of tasks because certain features are applicable to many different scenarios. For example, taking into account general measures of task complexity, past annotator performance, time on task, or comments provided by annotators can be informative irrespective of the underlying labeling task. Past research has rarely applied this approach [10] since most studies are based on archival data for which no annotator data (e.g., past performance) or task-completion data (e.g., time on task) are available.
1.2 Current Research
The current research presents a study on behavioral error detection models and their applications for search relevance annotation tasks. These tasks were drawn from an extensive data annotation program that generates training and evaluation data for search and recommendation systems. Here, we focus on search relevance annotations related to music streaming, video streaming, and mobile applications. A detailed description of the annotation task can be found in the Method section below.
Drawing on authentic data from the search relevance annotation program, we first use model-based error detection techniques to show that annotation errors can be detected with moderate classification performance. We demonstrate the predictability of errors for different search relevance rating tasks for all three ML applications: music streaming, video streaming, and mobile applications. Second, using Shapley values [20], we explore the contributions of a variety of annotator and task-related features in order to better understand which sources of information are driving the model’s predictions. Third, we explore the generalizability of our error detection models across rating tasks and ML applications. Fourth, we show that our error detection models can unlock considerable efficiency gains for auditing and relabeling by prioritizing tasks with high predicted error probabilities at the expense of tasks with low predicted error probabilities. Taken together, our findings show that error detection models can be highly effective in improving the efficiency of data annotation processes.
2 Method
2.1 Data Collection and Sampling
Our study is based on authentic data from an industry-scale search relevance annotation program. We used a large archival dataset of audited search relevance annotation tasks, including metadata, focused on music streaming, video streaming, and mobile applications. Audited tasks include annotator labels as well as ground truth labels from subject matter experts who were involved in the development of the annotation guidelines. We restricted the data to audited tasks in order to calculate error rates against ground truth labels provided by the subject matter experts. The dataset contained search relevance annotations for 42,611 individual tasks from 751 annotators spanning 23 storefronts (countries). From this dataset, 30% of observations were randomly selected for a holdout test set.
2.2 Search Relevance Annotation Tasks
The search relevance annotation tasks we used were presented as query-output pairs for which a human annotator provided a rating in one of 5 categories: Perfect, Excellent, Good, Acceptable, or Unacceptable. Each of the label categories was defined according to a set of rules and examples specified in comprehensive, task-specific guidelines provided to the annotators. In the case of music streaming, for example, returning the eponymous Metallica song for the query “Master of Puppets” would be rated as “Perfect”, while a Taylor Swift song would be rated as “Unacceptable”. In cases where none of the labels would apply, annotators could flag a problem in the annotation task. The current study covers a total of three different search relevance rating tasks for music streaming, video streaming, and mobile applications. Importantly, while each query would typically prompt a range of outputs or suggestions in each product’s user interface (UI), for the purpose of search relevance annotations, the individual outputs were rated in isolation. In other words, annotators only saw one query-output pair at a time and were blind to the ranking of the respective output item in the overall list of search results. Tasks were audited by subject matter experts who could overturn the original annotator labels.
p6cmp9.7cm Feature Name Description
Task Features
   input_occurrences Cumulative number of searches for focal query (reflects popularity)
   input_conversion_rate Proportion of users who converted on any output
   in_out_spacy_distance Cosine similarity between word embeddings of input and output string
   in_out_edit_distance Edit distance between input and output string
   input_media_type Indicates whether query was made through keyboard or verbally through a voice assistant
   output_media_type Type of search results in six different content categories
   input_misspelled Binary variable indicating if the input was misspelled
   input_language Categorical variable distinguishing between 21 different input languages
   input_query_type Intent classification for query
Past Performance Features
   error_N Rolling window of N = [7, 14, 21, 28] days for error rate of focal annotator (4 features)
   error_N_all Rolling window of N = [7, 14, 21, 28] days for error rate of all annotators in same eval (4 features)
   maj_error_N Rolling window of N = [7, 14, 21, 28] days for major error (annotator label more than 2 levels off) rate of focal annotator (4 features)
   maj_error_N_all Rolling window of N = [7, 14, 21, 28] days for major error (annotator label more than 2 levels off) rate for all annotators in same eval (4 features)
   error_N_diff Difference between annotator error rate and overall error rate of N = [7, 14, 21, 28] day rolling windows (4 features)
   vol_last_N Rolling window of N = [7, 14, 21, 28] days for task volume of focal annotator
   tenure_full_days Number of days a rater has been active since last deactivation
   tenure_updated_days Number of days a rater has been active since first joining the rater pool
   qualification_trials Number of trials an annotator took to pass the qualification test
   qualification_agreement_rate Performance on qualification test
   error_rolling_output_media_type_N Share of correct annotations by focal annotator for given media type over window of N = [1, 3, 5] (3 features)
   error_rolling_input_query_type_user_N Share of correct annotations by focal annotator for given input query type over window of N = [1, 3, 5] (3 features)
Session Context Features
   nth_task_in_session Counter indicating the position of the current task within an ongoing session
   seconds_into_session Time that has passed since the start of the session in seconds
Task-Completion Features
   answer_value Label provided by annotator
   time_on_task Time (in seconds) spent on focal annotation task
   comment_length Number of words in annotator comment
Table 1: Overview of features used in error models. Please refer to this table when interpreting the feature importance scores in Figure 2.